OpenStudy (anonymous):

if g(n) =n^2+n and h(n)=n^2-n which of the following is equal to h(m+1)? g(m), g(m)+1, g(m)-1, h(m)-1

OpenStudy (asnaseer):

first calculate what h(m+1) evaluates to

OpenStudy (anonymous):

m^2+2m+1-m-1 = ,m^2+m

OpenStudy (anonymous):

g(m)

OpenStudy (anonymous):

why

OpenStudy (anonymous):

@wrik003 why is it g(m

OpenStudy (anonymous):

because h(m+1) = ((m+1)^2) - (m+1) = (m^2) + m = g(m) (given)... and anyways I don't copy answers without doing the sums! ;-)